If James is 3 times the age of Eric, and 5 years ago James was 4 times the age of Eric how old are they each?: If James is 3 times the age of Eric, and 5 years ago James was 4 times the age of Eric how old are they each?
Answer by mangopeeler07(445)

First set variables for James and Eric. Since James is 3 times the age of Eric, and 5 years ago James was 4 times the age of Eric, you would set up two equations. Let's say that Eric is e. Then your equations would be james=3e and james-5=4(e-5). You can use the first equation james=3e and plug in that value for james in the second equation. Then it becomes 3e-5=4(e-5). Then distribute on the right side 3e-5=4e-20. Add five to both sides and get 3e=4e-15. Then subtract 4e from both sides and get -e=-15. cancel out the negatives and you are left with e=15.

Eric=e and James=3e. Plug in 15 for e and you get Eric=15 and James=45.
To check, subtract 5 from each and get 10 and 40. Since 40 is 10(4), then our ages are correct. Eric=15 and James=45.